OpenStudy (anonymous):

A bag of marbles had 16 blue marbles and the rest were green. For a game,5/8 of the 16 blue marbles were chosen, and 12 of the green marbles were also chosen.

OpenStudy (anonymous):

(a) How many of the bag's blue marbles were chosen? I thought 10 because I did 5/8*16=10 (b)How many of the bag's green marbles were not chosen? Not enough information (c) How many of the bag's marbles were chosen? 17.5 5/8*12=15/2=7.5+10=17.5

OpenStudy (anonymous):

@tester97

OpenStudy (tester97):

I agree with everything except Part c. If it is asking how many marbles were chosen you would just add the green and blue and it would equal 22 because \[\frac{ 5 }{ 8 }\times 16=10~then ~10+12=22\]

OpenStudy (anonymous):

oh... I wasn't supposed to take 5/8th of 12 :P thanks for checking!

OpenStudy (tester97):

Welcome :)